Find the focus and equation of the latus rectum if x^2+ky=0, where k>0
1
Expert's answer
2020-12-17T07:28:39-0500

Comparing the given equation "x^2=-ky" of parabola with standard equation "x^2=-4ay" , we obtain "k=4a"

"\implies a = k/4"


Focus of standard parabola is "(0,-a)"

So, for the given parabola, focus is "(0,-\frac{k}{4})"



Equation of latus rectum is given by "y=-a"


Hence for the given parabola, equation of latus rectum is, "y=-\frac{k}{4}"